Find the.equation of the line that passes through the points -3, -14 and 1 , -2
Found 2 solutions by swincher4391, lwsshak3:
Answer by swincher4391(1107)

(-3,-14) and (1,-2) has slope (-2+14)/(1-(-3) = 12 /4 = 3
y+14 = 3(x+3) <--- point slope form.
y +14 = 3x+9
y = 3x - 5 <--- y-intercept form
3x - y = 5 <--- standard form

Answer by lwsshak3(11628)

Find the.equation of the line that passes through the points -3, -14 and 1 , -2
***
Standard form for equation of a line: y=mx+b, m=slope, b=y-intercept
Using given points:
slope=∆y/∆x=(-2-(-14))/(1-(-3))=(-2+14)/(1+3)=12/4=3
equation: y=3x+b
solve for b using one of given points, say, (1,-2)
-2=3*1+b
b=-5
final equation: y=3x-5